40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6600     Entry time: Thu May 3 21:13:48 2012
Author: Koji 
Type: Summary 
Category: SUS 
Subject: ITMX/PRM/BS OPLEV aligned 

[Jenne/Den/Koji]

We locked Xarm/Yarm and manually alignmed ITMX/ITMY/BS/ETMX/ETMY/PZT1/PZT2.

ITMY OPLEV was largely misaligned ==> The beam was centered on the QPD.

----

Then we aligned PRM using SB locking PRMI.

We noticed that OPLEV servo does not work. It made the PRM just noiser.

We went into the PRM table and found that the OPLEV beam was clipped in the vacuum chamber.
We tried to maximize the reflected beam from the window by touching the steering mirrors at the injection side.

Then the reflected beam was introduced to the center of the QPD.

After the alignment, the OPLEV QPD SUM increased to 4000ish from 200ish.
According to the OPLEV trend data, this is a nominal value of the QPD SUM.

Now the OPLEV servo does not go crazy.

 --

BS OPLEV beam was centered on the QPD.

ELOG V3.1.3-